OpenStudy (anonymous):

Determine whether the graph of y = x2 + 2x − 8 has a maximum or minimum point, then find the maximum or minimum value. Maximum; (-1, -9) Minimum; (-1, -9) Maximum; (-9, -1) Minimum; (-9, -1)

OpenStudy (johnweldon1993):

Is this a parabola that opens UP or DOWN...?

OpenStudy (anonymous):

I think it opens up I"m not sure though.

OpenStudy (johnweldon1993):

You are correct...as long as that leading number (1x^2) is positive *as it is* the parabola opens up...and SINCE it opens up...  We have a minimum point right?

OpenStudy (anonymous):

yes.

OpenStudy (johnweldon1993):

Now we want to find the vertex of the parabola....so we will use the equation \[\large Vertex = \frac{ -b }{ 2a }\] Remember your equation is in the form \[\large ax^2 + bx + c\] So your 'b' value is 2....and your 'a' value is just 1 right?...so we have \[\large Vertex = \frac{ -2 }{ 2(1) }\] \[\large Vertex = \frac{ -2 }{ 2 }\] \[\large Vertex = -1 \] This is your 'x' coordinate of your vertex....already we can tell your answer choice is...?

OpenStudy (anonymous):

so the answer is B minimum (-1,-9)

OpenStudy (johnweldon1993):

That would be correct! And just for future reference...if you NEEDED to find the y-value of the vertex....you would simply plug in the x-value into your equation and solve for 'y'

OpenStudy (anonymous):

oh I knew about the plugin on how to find the y, so thank you for the help your a life savior.

OpenStudy (johnweldon1993):

No problem! :)